# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v4

@@
Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models
@@
Linear structural vector autoregressive models can be identified statistically without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian. We show that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes. We also discuss the problem of labelling the shocks, estimation of the parameters, and stationarity the model. The introduced methods are implemented to the accompanying R package sstvars. Our empirical application studies the effects of the climate policy uncertainty shock on the U.S. macroeconomy. In a structural logistic smooth transition vector autoregressive model with two regimes, we find that a positive climate policy uncertainty shock decreases production and increases inflation in times of both low and high economic policy uncertainty, but its inflationary effects are stronger in the periods of high economic policy uncertainty.
Keywords: smooth transition vector autoregression, threshold VAR, nonlinear SVAR, identification by non-Gaussianity, non-Gaussian SVAR
The author thanks Markku Lanne for the helpful discussions and comments as well as the Research Council of Finland for the financial support (Grant 347986).Contact address: Savi Virolainen, Faculty of Social Sciences, University of Helsinki, P. O. Box 17, FI–00014 University of Helsinki, Finland; e-mail: savi.virolainen@helsinki.fi. ORCiD ID: $0$-0002-5075-6821.
@@ 1 Introduction
Linear vector autoregressive (VAR) models are a standard tool in empirical macroeconomics due to their ability to effectively capture constant dynamic relationships of the included variables. Macroeconomic systems may, however, exhibit variation in their dynamics, induced by wars, crises, policy shifts, or business cycle fluctuations, for example. Such structural changes cannot be accommodated by linear VAR models, and therefore, nonlinear VAR models that are able capture such dynamics are often employed. Nonlinear structural VAR (SVAR) models are particularly useful, as they facilitate tracing out the causal effects of economic shocks, which may vary in time depending on the initial state of the economy as well as on the sign and size of the shock (Kilian and Lütkepohl, 2017, Chapter 18). To estimate these effects, the economic shocks need to be identified. Conventional identification methods typically employed in nonlinear SVAR models rely on restrictive assumptions, such as zero contemporaneous interactions among some of the variables (see Kilian and Lütkepohl, 2017, Section 18.2.2). Such restrictions with a clear interpretation are advantageous when based on economic reasoning, but often they are economically implausible and imposed just to achieve sufficient identification. To overcome the latter issue, statistical identification methods relying on the statistical properties of the data can be used.
@@
There are two main branches in the statistical identification literature: identification by heteroscedasticity (Rigobon, 2003, Lanne et al., 2010, Bacchiocchi and Fanelli, 2015, Lütkepohl and Netšunajev, 2017, Lewis, 2021, Virolainen, 2025b, and others) and identification by non-Gaussianity (Lanne et al., 2017, Lanne and Luoto, 2021, Lanne et al., 2023, and others). To the best our knowledge, this paper is the first to study identification by non-Gaussianity in nonlinear SVAR models. Under certain statistical conditions, both types of identification methods typically identify the shocks of a linear structural vector autoregressive (SVAR) model without imposing further restrictions. However, identification by heteroscedasticity without additional restrictions (as described in the cited literature) has the major drawback in nonlinear SVAR models that, for each shock, it restricts the relative magnitudes of the impact responses of the variables to stay constant over time (see, e.g., Lütkepohl and Netšunajev, 2017). This is an undesirable property because it is generally preferable to accommodate time-variation in the (relative) impact responses.
@@
This paper contributes to the literature on identification by non-Gaussianity by extending the framework proposed by Lanne et al. (2017) to smooth transition vector autoregressive (STVAR) models (e.g., Anderson and Vahid, 1998), which is major class of nonlinear SVAR models. We also consider the threshold vector autoregressive model of Tsay (1998) as a special case of the STVAR model. The STVAR model can flexibly capture nonlinear data generating dynamics by accommodating multiple regimes and gradual shifts between them, governed by the transition weighs. In contrast to its linear counterpart, the impact matrix of the structural STVAR model should generally allow for time-variation in the impact responses of the variables to the shocks, which complicates identification. Nonetheless, similarly to Lanne et al. (2017), it turns out that identification is achieved when the shocks are mutually independent and at most one of them is Gaussian. We show that under this condition, when the impact matrix of the STVAR model is defined as a weighted sum of the impact matrices of the regimes, the shocks are readily identified up to ordering and signs. The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be extended to other suitable weight functions as well.
@@
In line with the statistical identification literature, external information is required to label the identified structural shocks as economic shocks. Our nonlinear setup has the additional complication that the same shock should be assigned to the same column of the impact matrix across all regimes. Therefore, since our experience shows that the identification can be weak with respect to the ordering and signs of the columns of the regime-specific impact matrices, we recommend adopting a blended identification strategy that combines identification by non-Gaussianity with additional identifying information (cf. Carriero et al., 2024).
Developments related to ours include the new framework of Morioka et al. (2021) called independent innovation analysis (IIA). The IIA estimates the (conditionally) contemporaneously independent innovations of a nonlinear SVAR model without imposing any specific functional form. Under fairly general assumptions, their method consistently estimates the independent innovations up to permutation and scalar component-wise invertible transformations. While the IIA does seem promising, it is not obvious how to apply it in empirical macroeconometrics, as identification only up to such transformations is, in our understanding, insufficient for structural analysis. Moreover, Morioka et al. (2021) assume the innovations follow a distribution in the exponential family, thereby excluding various interesting alternatives. In contrast, our approach facilitates estimation of the parameters of an identified nonlinear SVAR model, enabling structural analysis without restricting the distribution of independent shocks beyond ruling out Gaussian and infinite-variance cases.
@@
Our empirical application studies the macroeconomic effects of climate policy uncertainty (shocks) and considers monthly U.S. data from 1987:4 to 2024:12. Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure climate policy uncertainty (CPU) with the CPU index (Gavriilidis, 2021), which is constructed based on the amount of newspaper coverage on topics related to CPU. We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty (EPU). Therefore, we fit a two-regime structural logistic STVAR model using the first lag of the EPU index (Baker et al., 2016) as the switching variable. We find that a positive CPU shock decreases production and increases inflation in times of both low and high EPU, but its inflationary effects are stronger in the periods of high EPU. Our results are, hence, in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).
@@
The rest of this paper is organized as follows. Section 2 presents the framework of reduced form STVAR models and provides examples of the covered models. Section 3 discusses identification of the shocks in structural STVAR models, presents our results on identification by non-Gaussianity, and discusses the problem of labelling the shocks. Section 4 discusses stationarity of the model and proposes estimating its parameters with a penalized likelihood -based estimator using a three-step procedure. Section 5 presents the empirical application and Section 6 concludes. Appendices provide detailed discussion on ergodic stationarity of the model; proofs for the stated lemmas, propositions, and theorem; details related to the proposed estimation procedure; a Monte Carlo study assessing the performance of the proposed estimator; and details related to the empirical application. Finally, the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
@@ 2 Reduced form STVAR models
Let $y_{t}$, $t=1,2,...$, be the $d$-dimensional time series of interest and $\mathcal{F}_{t-1}$ denote the $\sigma$-algebra generated by the random vectors $\{y_{t-j},j>0\}$. We consider STVAR models with $M$ regimes and autoregressive order $p$ assumed to satisfy
@@
where $\phi_{1},...,\phi_{M}\in\mathbb{R}^{d}$ are the intercept parameters; $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, $i=1,...,p$, are the autoregression matrices; $u_{t}$ is a martingale difference sequence of reduced form innovations; $\Omega_{y,t}$ is the positive definite conditional covariance matrix of $u_{t}$ (conditional on $\mathcal{F}_{t-1}$), which may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$; and $\nu$ collects any other parameters that the distribution of $u_{t}$ depends on. The transition weights $\alpha_{m,t}$ are assumed to be either exogenous (nonrandom) or $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, and to satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$. Thus, the types of transition weights typically used in macroeconomic applications, including exogenous, logistic, and threshold weights discussed below, are accommodated. The transition weights express the proportions of the regimes the process is on at each point of time and determine how the process shifts between them.
It is easy to see that, conditional on $\mathcal{F}_{t-1}$, the conditional mean of the above-described process is $\mu_{y,t}\equiv E[y_{t}|\mathcal{F}_{t-1}]=\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}$, a weighted sum the regime-specific means $\mu_{m,t}$ with the weights given by the transition weights $\alpha_{m,t}$. The conditional covariance matrix is $\Omega_{y,t}$, but discussion on its specification is postponed to Section 3. See Hubrich and Teräsvirta (2013) for a survey on STVAR literature, including also specifications more general than ours.
Our framework accommodates various types of transition weights employed in empirical macroeconomics. To provide a few examples, consider first the threshold VAR (TVAR) model (Tsay, 1998), which is obtained from (2.1) and (2.2) by assuming that the transition weights are defined as
$\alpha_{m,t}=\left\{\begin{matrix}1&\text{if}\ \ r_{m-1}<z_{t}\leq r_{m},\\ 0&\text{otherwise},\phantom{aaaaaaa}\end{matrix}\right.$ (2.3)
where $r_{0}\equiv-\infty$, $r_{M}\equiv\infty$, $m=1,...,M$, are threshold parameters, and $z_{t}$ is a switching variable. At each $t$‚ the model defined in Equations (2.1), (2.2) and (2.3) reduces to a linear VAR corresponding to one of the regimes that is determined according to the level of the switching variable $z_{t}$, which can be an exogenous variable, a lagged endogenous variable, or a linear combination of multiple such variables, for example.
Another popular specification is to assume two regimes ($M=2$) and logistic transition weights. In the logistic STVAR model (Anderson and Vahid, 1998), the transition weights vary according to a logistic function as
@@
where $z_{t}$ is the switching variable, $c\in\mathbb{R}$ is a location parameter, and $\gamma>0$ is a scale parameter. The location parameter $c$ determines the midpoint of the transition function, i.e., the value of the switching variable when the weights are equal. The scale parameter $\gamma$, in turn, determines the smoothness of the transitions (smaller $\gamma$ implies smoother transitions), and it is assumed strictly positive so that $\alpha_{2,t}$ is increasing in $z_{t}$.
In addition to specifying the transition weights or the error process $u_{t}$ in various ways, different types of models can be obtained by imposing parameter restrictions to the model defined in (2.1)-(2.2). For instance, a model that is linear in the autoregressive (AR) coefficients is obtained by assuming that they are constant across the regimes, i.e., $A_{m,i}=A_{i}$ and $\phi_{m}=\phi$ for all $i=1,...,p$ and $M=1,...,M$. The model then simplifies to
$y_{t}=\phi+\sum_{i=1}^{p}A_{i}y_{t-i}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu).$ (2.5)
The linear model in (2.5) is typically substantially easier to estimate and has much fewer parameters than the fully nonlinear specification in (2.1)-(2.2), making it more feasible for large scale SVAR analyses, but it can still accommodate shifts in the volatility regime through the error process $u_{t}$.
@@ 3.1 Structural STVAR models
A structural STVAR model is obtained from the reduced form model defined in Section 2 by identifying the serially and mutually uncorrelated structural shocks $e_{t}=(e_{1t},...,e_{dt})$ $(d\times 1)$ from the reduced form innovations $u_{t}$. Specifically, the structural shocks are recovered from the reduced form shocks with the transformation
@@
where $B_{y,t}$ is an invertible ($d\times d$) impact matrix that governs the contemporaneous relationships of the shocks and may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$. In other words, assuming a unit variance normalization for the structural shocks, the identification problem amounts to finding an impact matrix $B_{y,t}$ such that $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$. The impact matrix is not, however, generally uniquely identified without imposing further restrictions on the model, since there exists many such decompositions of $\Omega_{y,t}$. Various types of restrictions that recover the structural shocks have been proposed in the literature, many of which are discussed in Kilian and Lütkepohl (2017), for example.
Following the literature on statistical identification by non-Gaussianity in the framework proposed by Lanne et al. (2017), our identifying restriction is that the shocks are mutually independent and at most one of them is Gaussian. Under this assumption, unlike in the Gaussian case, different decompositions of the conditional covariance matrix $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$ generally lead to observationally distinct models, effectively identifying a static impact matrix and the shocks (Lanne et al., 2017). We extend this result to a time-varying impact matrix, and therefore, it is useful to parametrize the structural model directly with $B_{y,t}$. Hence, we make the identity $u_{t}=B_{y,t}e_{t}$ explicit in Equation (2.1) as
@@
To incorporate time-variation in the impact matrix, we specify its functional form. A natural specification for STVAR models is to assume a constant impact matrix for each of the regimes and define the impact matrix of the process to be their weighted sum as
@@
where $B_{1},...,B_{M}$ are invertible $(d\times d)$ impact matrices of the regimes. Also the impact matrix $B_{y,t}$ needs to be invertible to ensure positive definiteness of the conditional covariance matrix, which does not automatically follow from the invertibility of $B_{1},...,B_{M}$. Nevertheless, it turns out that $B_{y,t}$ defined in (3.3) is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, as is stated in the following lemma (which is proven in Appendix B.1).
@@
Suppose $B_{1},...,B_{M}$ are invertible $(d\times d)$ matrices and $\alpha_{1,t},...,\alpha_{M,t}$ scalars such that $\sum_{m=1}^{M}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$ for all $t$ and $m=1,...,M$. Then, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
The result of Lemma 1 holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that the set where the matrices $B_{1},...,B_{M}$ are such that $B_{y,t}$ is singular for some $t$ has a Lebesgue measure zero. This implies that invertible matrices $B_{1},...,B_{M}$ generally lead to an invertible impact matrix $B_{y,t}$, excluding some special cases of $B_{1},...,B_{M}$ for which this result does not hold (e.g., if $B_{2}=-B_{1}$, then $B_{y,t}=0$ when $\alpha_{1,t}=\alpha_{2,t}=0.5$). To further improve intuition, note that the result of Lemma 1 implies that if the matrices $B_{1},...,B_{M}$ are drawn by random from a continuous distribution, the probability of obtaining matrices such that $B_{y,t}$ is singular for some $t$ is zero. In practice, the parameter space should be simply restricted in estimation so that parameter values leading to $B_{y,t}$ that is singular for some $t$ are excluded.
Assuming impact matrix of the form (3.3), the conditional covariance matrix of $y_{t}$, conditional on $\mathcal{F}_{t-1}$, is obtained as
@@
where $\Omega_{m}\equiv B_{m}B_{m}^{\prime}$ and $\Omega_{m,n}\equiv B_{m}B_{n}^{\prime}$. The conditional covariance matrix of $y_{t}$ can be thus described as a weighted sum of $M^{2}$ matrices with the weights varying in time according to the transition weights $\alpha_{m,t}$, $m=1,...,M$. If the process is completely in one of the regimes, i.e., $\alpha_{m,t}=1$ for some $m\in\{1,...,M\}$, then $\Omega_{y,t}$ reduces to $\Omega_{m}$, implying that this constitutes as the conditional covariance matrix of Regime $m$. When the process is not completely in any of the regimes, $\Omega_{y,t}$ depends on both the regime-specific covariance matrices and the cross terms. Hence, this specification of the structural model is different to the conventional reduced form STVAR specification in which the conditional covariance matrix of $y_{t}$ is a weighted sum of the covariance matrices of the regimes. Nonetheless, our model specification facilitates exploiting non-Gaussianity of the shocks in their identification, as we show in the next section.

### v5

# Identification by non-Gaussianity in structural smooth transition vector autoregressive models

@@
Identification by non-Gaussianity in structural smooth transition vector autoregressive models
@@
We show that structural smooth transition vector autoregressive models are statistically identified if the shocks are mutually independent and at most one of them is Gaussian. This extends a known identification result for linear structural vector autoregressions to a time-varying impact matrix. We also propose an estimation method, show how a blended identification strategy can be adopted to address weak identification, and establish a sufficient condition for ergodic stationarity. The introduced methods are implemented in the accompanying R package sstvars. Our empirical application finds that a positive climate policy uncertainty shock reduces production and raises inflation under both low and high economic policy uncertainty, but its effects, particularly on inflation, are stronger during the latter.
Keywords: smooth transition vector autoregression, nonlinear SVAR, identification by non-Gaussianity, non-Gaussian SVAR, climate policy uncertainty
The author thanks Markku Lanne for the helpful discussions and comments as well as the Research Council of Finland for the financial support (Grant 347986).Contact address: Savi Virolainen, Faculty of Social Sciences, University of Helsinki, P. O. Box 17, FI–00014 University of Helsinki, Finland; e-mail: savi.virolainen@helsinki.fi. ORCiD ID: 0000-0002-5075-6821.The author does not have conflicts of interest to declare.
@@ 1 Introduction
Linear structural vector autoregressive (SVAR) models have become a standard tool in empirical macroeconomics, as they can effectively capture constant dynamic relationships among the variables and facilitate tracing out the causal effects of economic shocks. Macroeconomic systems may, however, exhibit variation in their dynamics, induced by crises, policy shifts, or business cycle fluctuations, for example. Hence, also the effects of the shocks may depend on the state of the economy. Such features cannot be accommodated by linear SVAR models, and therefore nonlinear SVAR models that are able to capture such features are often employed (see, e.g., Kilian and Lütkepohl, 2017, Chapter 18). In nonlinear SVAR models, the structural shocks are typically identified with conventional methods that impose economically interpretable but restrictive assumptions, such as zero contemporaneous interactions among some of the variables, which are not always plausible in practice. To address this issue, statistical identification methods relying on the statistical properties of the data can be used (see Kilian and Lütkepohl, 2017, Chapter 14).
@@
There are two main branches in the statistical identification literature: identification by heteroskedasticity (Rigobon, 2003, Lanne and Lütkepohl, 2010, Bacchiocchi and Fanelli, 2015, Lütkepohl and Netšunajev, 2017, Lewis, 2021, Virolainen, 2025b, and others) and identification by non-Gaussianity (Lanne et al., 2017, Lanne and Luoto, 2021, Lanne et al., 2023, and others). To the best of our knowledge, this paper is the first to examine identification by non-Gaussianity in nonlinear SVAR models. Under certain statistical conditions, both types of statistical information typically suffice to identify the shocks in a linear SVAR model. As pointed out by Lütkepohl and Netšunajev (2017), among others, identification by heteroskedasticity without additional restrictions, however, has the major drawback in nonlinear SVAR models that it imposes time-invariant (relative) impact effects of the shocks. We show that identification by non-Gaussianity, in turn, facilitates recovering the shocks without such undesirable restrictions.
@@
This paper contributes to the literature on identification by non-Gaussianity by extending the framework of Lanne et al. (2017) to structural smooth transition vector autoregressive (STVAR) models, which is a major class of nonlinear SVAR models (see, e.g., Hubrich and Teräsvirta, 2013). The STVAR model can flexibly capture nonlinear data generating dynamics by accommodating multiple regimes and gradual as well as abrupt shifts between them, governed by the transition weights. In contrast to its linear counterpart, the impact matrix of the structural STVAR model should generally allow for time-variation in the impact responses of the variables to the shocks, which complicates identification. Nevertheless, similarly to Lanne et al. (2017), it turns out that identification is achieved when the shocks are mutually independent and at most one of them is Gaussian. We show that under this condition, the shocks are readily identified up to ordering and signs when the impact matrix of the STVAR model is defined as a weighted sum of the impact matrices of the regimes. While we focus on exogenous, logistic, and threshold transition weights, our results can be extended to other suitable weight functions as well.
@@
In line with the statistical identification literature, external information is required to label the identified structural shocks as economic shocks. Our nonlinear setup has the additional complication that the same shock should be assigned to the same column of the impact matrix across all regimes. Moreover, our experience shows that identification is often weak in the sense that there are often multiple local solutions with a fit close to the global solution. We find such local solutions to be often largely quite similar to each other, with the differences frequently but not exclusively related to different ordering and signs of the columns of the regime-specific impact matrices. Nonetheless, since different local solutions may produce different results in structural analysis, we recommend addressing weak identification by adopting a blended identification strategy that combines identification by non-Gaussianity with supplementary identifying information (cf. Carriero et al., 2024).
@@
Our methods are demonstrated in an empirical application to the macroeconomic effects of climate policy uncertainty (shocks) that considers monthly U.S. data from 1987:4 to 2024:12. Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure climate policy uncertainty (CPU) with the CPU index of Gavriilidis (2021), which is constructed based on the amount of newspaper coverage on topics related to CPU. We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty (EPU). Therefore, we fit a two-regime structural logistic STVAR model using the first lag of the EPU index (Baker et al., 2016) as the switching variable. We find that a positive CPU shock reduces production and raises inflation in times of both low and high EPU, but its effects, particularly on inflation, are stronger in the periods of high EPU. Our results are, hence, in line with the previous literature suggesting that a positive CPU shock reduces production and raises inflation (Khalil and Strobel, 2023 and Huang and Punzi, 2024), while Fried et al. (2022) found that it decreases production.
@@
The rest of this paper is organized as follows. Section 2 presents the framework of reduced form STVAR models. Section 3 discusses identification of the shocks in structural STVAR models, presents our identification results, and discusses the problem of labeling the shocks as well as how to address weak identification by adopting a blended identification strategy. Section 4 discusses stationarity of the model and proposes estimating its parameters with a penalized likelihood-based estimator using a three-step procedure. Section 5 presents the empirical application and Section 6 concludes. Further details can be found in the appendices, and the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
@@ 2 Reduced form STVAR models
Let $y_{t}$, $t=1,2,$..., be the $d$-dimensional time series of interest and $\mathcal{F}_{t-1}$ denote the $\sigma$-algebra generated by the random vectors $\{y_{t-j},j>0\}$ (i.e., $\mathcal{F}_{t-1}$ contains the information about the history of the process). We consider STVAR models with $M$ regimes and autoregressive order $p$ assumed to satisfy
@@
where $\phi_{1},...,\phi_{M}\in\mathbb{R}^{d}$ are the intercept parameters; $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, i=1,,pi=1,...,p, are the autoregression matrices; $u_{t}$ is a martingale difference sequence of reduced form innovations; $\Omega_{y,t}$ is the positive definite conditional covariance matrix of $u_{t}$ (conditional on $\mathcal{F}_{t-1}$), which may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$; and $\nu$ collects any other parameters that the distribution of $u_{t}$ depends on.
The transition weights $\alpha_{m,t}$ are assumed to be either exogenous (nonrandom) or $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, and to satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$. Thus, we accommodate the types of transition weights typically used in macroeconomic applications, including exogenous, logistic, and threshold weights discussed below. The transition weights express the proportions of the regimes the process is in at each point of time and determine how the process shifts between them.
It is easy to see that, conditional on $\mathcal{F}_{t-1}$, the conditional mean of the above-described process is $\mu_{y,t}\equiv E[y_{t}|\mathcal{F}_{t-1}]=\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}$, a weighted sum the regime-specific means $\mu_{m,t}$ with the weights given by the transition weights $\alpha_{m,t}$. The discussion on the specification of the conditional covariance matrix $\Omega_{y,t}$ is postponed to Section 3. The linear vector autoregressive (VAR) model is obtained as a special case by assuming constant autoregressive dynamics across the regimes, i.e., $A_{m,i}=A_{i}$ and $\phi_{m}=\phi$ for all i=1,,pi=1,...,p and m=1,,Mm=1,...,M. [^1] See Hubrich and Teräsvirta (2013) for a survey on STVAR literature, including also specifications more general than ours.
A popular specification is obtained by assuming two regimes ($M=2$) and logistic transition weights. In the logistic STVAR model (Anderson and Vahid, 1998), the transition weights vary according to a logistic function as
@@
where $c\in\mathbb{R}$ is the location parameter, $\gamma>0$ is the scale parameter, and $z_{t}$ is the switching variable, which we assume to be a lagged endogenous variable up to the lag $p$ (that is, $z_{t}\in\{y_{it-j},i=1,...,d,j=1,...,p\}$). The location parameter $c$ determines the midpoint of the transition function, i.e., the value of the switching variable when the weights are equal. The scale parameter $\gamma$, in turn, determines the smoothness of the transitions (the smaller $\gamma$ is, the smoother the transition is), and it is assumed strictly positive so that $\alpha_{2,t}$ is increasing in $z_{t}$. In the special case when $\gamma\rightarrow\infty$, regime-switches are discrete and the logistic weights reduce to the threshold weights of Tsay (1998) (see Appendix A for details).
@@ 3.1 Structural STVAR models
A structural STVAR model is obtained from the reduced form model defined in Section 2 by identifying the serially and mutually uncorrelated structural shocks $e_{t}=(e_{1t},...,e_{dt})$ $(d\times 1)$ from the reduced form innovations $u_{t}$. Specifically, the structural shocks are recovered from the reduced form innovations with the transformation
@@
where $B_{y,t}$ is an invertible ($d\times d$) impact matrix that governs the contemporaneous relationships of the shocks and may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$. In other words, assuming a unit variance normalization for the structural shocks, the identification problem amounts to finding an impact matrix $B_{y,t}$ such that the conditional covariance matrix $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$. However, $B_{y,t}$ is not generally uniquely identified without further restrictions, as many such decompositions exist. Various identification methods have been proposed to resolve this issue (see, e.g., Kilian and Lütkepohl, 2017).
Lanne et al. (2017), among others, have shown that if the shocks are mutually independent and at most one of them is Gaussian, a static impact matrix is readily identified (up to ordering and signs of its columns) without additional restrictions. We extend this result to a time-varying impact matrix, and therefore, it is useful to parametrize the structural model directly with $B_{y,t}$. Hence, we make the identity $u_{t}=B_{y,t}e_{t}$ explicit in Equation (2.1) as
@@
To incorporate time-variation in the impact matrix, we specify its functional form. A natural specification for structural STVAR models is to assume a constant impact matrix for each of the regimes and define the impact matrix of the process to be their weighted sum as
@@
where $B_{1},...,B_{M}$ are invertible $(d\times d)$ impact matrices of the regimes. Also the impact matrix $B_{y,t}$ needs to be invertible to ensure positive definiteness of the conditional covariance matrix, which does not automatically follow from the invertibility of $B_{1},...,B_{M}$. Nevertheless, it turns out that $B_{y,t}$ defined in (3.3) is invertible for all $t$ almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, as is stated in the following lemma (which is proven in Appendix C.1).
@@
Suppose $B_{1},...,B_{M}$ are $(d\times d)$ matrices and $\alpha_{1,t},...,\alpha_{M,t}$ scalars such that $\sum_{m=1}^{M}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$ for all $t$ and m=1,,Mm=1,...,M. Then, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$ almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
The result of Lemma 1 holds almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that the set where the matrices $B_{1},...,B_{M}$ are such that $B_{y,t}$ is singular for some $t$ has Lebesgue measure zero. In other words, invertible matrices $B_{1},...,B_{M}$ generally lead to an invertible impact matrix $B_{y,t}$, excluding some special cases of $B_{1},...,B_{M}$ for which this result does not hold (e.g., if $B_{2}=-B_{1}$, then $B_{y,t}=0$ when $\alpha_{1,t}=\alpha_{2,t}=0.5$). To support intuition, note that the result of Lemma 1 implies that if the matrices $B_{1},...,B_{M}$ are drawn by random from a continuous distribution, the probability of obtaining matrices such that $B_{y,t}$ is singular for some $t$ is zero.
Assuming an impact matrix of the form (3.3), the conditional covariance matrix of $y_{t}$, conditional on $\mathcal{F}_{t-1}$, is obtained as
@@
where $\Omega_{m}\equiv B_{m}B_{m}^{\prime}$ and $\Omega_{m,n}\equiv B_{m}B_{n}^{\prime}$. The conditional covariance matrix of $y_{t}$ can thus be described as a weighted sum of $M^{2}$ matrices with the weights varying in time according to the transition weights $\alpha_{m,t}$, m=1,,Mm=1,...,M. If the process is completely in one of the regimes, i.e., $\alpha_{m,t}=1$ for some $m\in\{1,...,M\}$, then $\Omega_{y,t}$ reduces to $\Omega_{m}$, implying that this constitutes the conditional covariance matrix of Regime $m$. When the process is not completely in any of the regimes, $\Omega_{y,t}$ depends on both the regime-specific covariance matrices and the cross terms. Hence, this specification of the structural model is different to the conventional reduced form STVAR specification in which the conditional covariance matrix of $y_{t}$ is a weighted sum of the covariance matrices of the regimes. Nevertheless, our model specification facilitates exploiting non-Gaussianity of the shocks in their identification as is shown in the next section.